# Algorithms for mean-field variational inference via polyhedral optimization in the Wasserstein space

### v1

@@ 6 Conclusion
As for our application to mean-field VI, one direction to explore would be to explicitly quantify the dependence of the parameter $\Upsilon$ on $\mathcal{M}$, thereby initiating the search for optimal choices of $\mathcal{M}$. We also ask if our analysis of SGD can be improved, e.g., by removing the boundedness assumption in $(\Xi)$. Finally, another important question is to move beyond the well-conditioned setting.
@@ Acknowledgements
The authors are grateful to Jonathan Niles-Weed for helpful discussions at all stages of this work. YJ acknowledges support from NYU through the SURE program, where this project originated under the supervision of AAP. SC acknowledges the support of the Eric and Wendy Schmidt Fund at the Institute for Advanced Study. AAP acknowledges the support of Meta AI research, as well as the National Science Foundation through NSF Award 1922658.
@@ Appendix A Proofs for Section 3
$\leq 2h^{2}\,\mathbb{E}[\mathbb{E}_{t}\|Q^{-1}\,(\hat{\nabla}_{\lambda}{\mathcal{F}}_{t}-\nabla_{\lambda}{\mathcal{F}}_{t})\|_{Q}^{2}]+2M^{2}h^{2}\,\mathbb{E}\|\lambda^{t}-\lambda^{\star}\|^{2}_{Q}\,.$
@@
$\mathbb{E}\|\lambda^{(t+1)}-\lambda^{\star}\|^{2}_{Q}$ $\leq(1-\tfrac{mh}{2})^{t}\,\mathbb{E}\|\lambda^{(t)}-\lambda^{\star}\|^{2}_{Q}+\frac{2c_{0}h}{m}$
$\leq\exp(-mht/2)\,\mathbb{E}\|\lambda^{(t)}-\lambda^{\star}\|^{2}_{Q}+\frac{2c_{0}h}{m}\,.$
@@ Appendix C Proofs for Section 5
$\log\pi_{i}^{\star}\circ T_{i}^{\star}(x)=-\frac{x^{2}}{2}-(\log T_{i}^{\star})^{\prime}(x)\,.$ (19)
@@
$(\log\pi^{\star}_{i}\circ T^{\star}_{i})^{\prime}(x)=V_{i}^{\prime}(T^{\star}_{i}(x))\,(T_{i}^{\star})^{\prime}(x)=-x-\frac{(T_{i}^{\star})^{\prime\prime}(x)}{(T_{i}^{\star})^{\prime}(x)}\,.$ (20)
@@
$(T_{i}^{\star})^{\prime\prime}(x)=-(T_{i}^{\star})^{\prime}(x)\,\bigl{(}x+V_{i}^{\prime}(T_{i}^{\star}(x))\,(T^{\star}_{i})^{\prime}(x)\bigr{)}\,.$ (21)
@@
$(\log\pi_{i}^{\star}\circ T_{i}^{\star})^{\prime\prime}(x)$ $=V_{i}^{\prime\prime}(T^{\star}_{i}(x))\,(T_{i}^{\star})^{\prime}(x)^{2}+V_{i}^{\prime}(T^{\star}_{i}(x))\,(T_{i}^{\star})^{\prime\prime}(x)$
@@
$(T_{i}^{\star})^{\prime\prime\prime}(x)$ $=\frac{(T_{i}^{\star})^{\prime\prime}(x)^{2}}{(T_{i}^{\star})^{\prime}(x)}-(T_{i}^{\star})^{\prime}(x)\,\bigl{(}1+V_{i}^{\prime\prime}(T_{i}^{\star}(x))\,(T_{i}^{\star})^{\prime}(x)^{2}+V_{i}^{\prime}(T_{i}^{\star}(x))\,(T_{i}^{\star})^{\prime\prime}(x)\bigr{)}\,.$
@@
$|(T_{i}^{\star})^{\prime\prime}(x)^{2}/(T_{i}^{\star})^{\prime}(x)|$ $\lesssim\frac{\kappa^{3/2}}{\sqrt{\ell_{V}}}\,(1+|x|^{2})\,,$
@@
Hence, the final bound scales as
@@
$\|\bar{T}-\hat{T}\|_{L^{2}(\rho)}^{2}\leq d\varepsilon_{0}^{2}\qquad\text{and}\qquad\|D(\bar{T}-\hat{T})\|_{L^{2}(\rho)}^{2}\leq d\varepsilon_{1}^{2}\,.$ (22)
@@
We first make a series of reductions. By assumption, $D\bar{T}\succeq\alpha I$, and by definition, $\hat{T}$ is of the form $\alpha\,{\operatorname{id}}+\sum_{T\in\mathcal{M}}\lambda_{T}T+v$. By replacing $\bar{T}$ with $\bar{T}-\alpha\operatorname{id}$, it suffices to prove the following statement: assuming that $0\preceq D\bar{T}\preceq\ell_{V}^{-1/2}\,I$ together with the second derivative bound on $\bar{T}$, there exists $\hat{T}$ of the form $\sum_{T\in\mathcal{M}}\lambda_{T}T+v$ such that (22) holds. However, from the structure of $\mathcal{M}$, the problem now separates across the coordinates and it suffices to prove this statement with $d=1$.
Truncation procedure. We will construct $\hat{T}$ so that $\bar{T}(-R)=\hat{T}(-R)$ and $\bar{T}(+R)=\hat{T}(+R)$. Assuming that this holds, the bound on $\bar{T}^{\prime}$ and the fact that $\hat{T}$ is constant on $(-\infty,-R]$ and on $[+R,+\infty)$ readily implies
$|\bar{T}(x)-\hat{T}(x)|\leq\frac{1}{2\sqrt{\ell_{V}}}\,(|x|-R)\,,\qquad\text{for}\leavevmode\nobreak\ |x|\geq R\,.$
@@
$\int_{\mathbb{R}\setminus(-R,+R)}|\bar{T}-\hat{T}|^{2}\,\mathrm{d}\rho\leq\frac{1}{4\sqrt{2\uppi}\,\ell_{V}}\int_{\mathbb{R}\setminus(-R,+R)}{(|x|-R)}^{2}\exp(-x^{2}/2)\,\mathrm{d}x\,.$
@@
Standard Gaussian tail bounds and the Cauchy–Schwarz inequality imply that with the choice $R\asymp\sqrt{\log(1/(\ell_{V}\varepsilon^{2}))}$, we obtain $\|\bar{T}-\hat{T}\|_{L^{2}(\rho)}^{2}\vee\|\bar{T}^{\prime}-\hat{T}^{\prime}\|_{L^{2}(\rho)}^{2}\lesssim\varepsilon^{2}$.
@@
Size of the generating family. Finally, the size of $\mathcal{M}$ is $O(R/\delta)=\widetilde{O}(\kappa^{1/2}d^{1/4}/\varepsilon^{1/2})$, which completes the proof. ∎
@@
The key property of the piecewise cubics is that $(\psi^{\rm cub,\pm})^{\prime}(0)=(\psi^{\rm cub,\pm})^{\prime}(1)=0$. This means that even after adding the piecewise cubics, $\bar{T}^{\prime}$ and $\hat{T}^{\prime}$ agree at all of the endpoints of the sub-intervals. However, we must check that adding these piecewise cubics does not destroy the approximation rates (24) and (25). Since $|(\psi^{\rm cub,\pm}_{I})^{\prime}|\lesssim 1/\delta$, the bound on the coefficients for the piecewise cubics shows that the derivative of the piecewise cubic part of $\hat{T}$ is bounded in magnitude by $O(\kappa^{2}R^{2}\delta^{2}/\sqrt{\ell_{V}})$, so that (24) is intact. Similarly, (25) is also intact, either by integrating (24) or by using the bound on the coefficients of the piecewise cubics. The proof is concluded via bookkeeping. ∎
@@
Assuming the first, we can prove the second by applying the Brascamp–Lieb inequality (Brascamp and Lieb,, 1976):

### v4

@@ 8 Conclusion
As for our application to mean-field VI, one direction to explore would be to see if our analysis of SGD can be improved, e.g., by removing the boundedness assumption in $(\Xi)$. Finally, another important question is to move beyond the well-conditioned setting.
@@ Acknowledgements
The authors are grateful to Jonathan Niles-Weed for helpful discussions at all stages of this work, as well as three reviewers that greatly improved the quality of our manuscript. YJ acknowledges support from NYU through the SURE program, where this project originated under the supervision of AAP. SC acknowledges the support of the Eric and Wendy Schmidt Fund at the Institute for Advanced Study. AAP acknowledges the support of Meta AI research, as well as the National Science Foundation through NSF Award 1922658.
@@ Appendix A Proofs for Section 3
$\leq 2h^{2}\,\mathbb{E}[\mathbb{E}_{t}\|Q^{-1}\,(\hat{\nabla}_{\lambda}{\mathcal{F}}_{t}-\nabla_{\lambda}{\mathcal{F}}_{t})\|_{Q}^{2}]+2M^{2}h^{2}\,\mathbb{E}\|\lambda^{(t)}-\lambda^{\star}\|^{2}_{Q}\,.$
@@
$\mathbb{E}\|\lambda^{({t})}-\lambda^{\star}\|^{2}_{Q}$ $\leq(1-\tfrac{mh}{2})^{t}\,\|\lambda^{(0)}-\lambda^{\star}\|^{2}_{Q}+\frac{2c_{0}h}{m}$
$\leq\exp(-mht/2)\,\|\lambda^{(0)}-\lambda^{\star}\|^{2}_{Q}+\frac{2c_{0}h}{m}\,.$
@@ Appendix C Proofs for Section 5
$\log\pi_{i}^{\star}\circ T_{i}^{\star}(x)=-\frac{x^{2}}{2}-\log(T_{i}^{\star})^{\prime}(x)-\tfrac{1}{2}\log(2\uppi)\,.$ (20)
@@
$(\log\pi^{\star}_{i}\circ T^{\star}_{i})^{\prime}(x)={-}V_{i}^{\prime}(T^{\star}_{i}(x))\,(T_{i}^{\star})^{\prime}(x)=-x-\frac{(T_{i}^{\star})^{\prime\prime}(x)}{(T_{i}^{\star})^{\prime}(x)}\,.$ (21)
@@
$(T_{i}^{\star})^{\prime\prime}(x)=-(T_{i}^{\star})^{\prime}(x)\,\bigl{(}x{-}V_{i}^{\prime}(T_{i}^{\star}(x))\,(T^{\star}_{i})^{\prime}(x)\bigr{)}\,.$ (22)
@@
$(\log\pi_{i}^{\star}\circ T_{i}^{\star})^{\prime\prime}(x)$ $={-}\bigl{(}V_{i}^{\prime\prime}(T^{\star}_{i}(x))\,(T_{i}^{\star})^{\prime}(x)^{2}+V_{i}^{\prime}(T^{\star}_{i}(x))\,(T_{i}^{\star})^{\prime\prime}(x)\bigr{)}$
@@
$(T_{i}^{\star})^{\prime\prime\prime}(x)$ $=\frac{(T_{i}^{\star})^{\prime\prime}(x)^{2}}{(T_{i}^{\star})^{\prime}(x)}-(T_{i}^{\star})^{\prime}(x)\,\bigl{(}1{-}V_{i}^{\prime\prime}(T_{i}^{\star}(x))\,(T_{i}^{\star})^{\prime}(x)^{2}{-}V_{i}^{\prime}(T_{i}^{\star}(x))\,(T_{i}^{\star})^{\prime\prime}(x)\bigr{)}\,.$
@@
$|(T_{i}^{\star})^{\prime\prime}(x)^{2}/(T_{i}^{\star})^{\prime}(x)|$ $\lesssim\frac{\kappa^{{2}}}{\sqrt{\ell_{V}}}\,(1+|x|^{2})\,,$
@@
To obtain the first bound, note that by (22) and the subsequent calculations, we have $|(T_{i}^{\star})^{\prime\prime}(x)/(T_{i}^{\star})^{\prime}(x)|\lesssim\kappa\,(1+|x|)$. Square and use $(T_{i}^{\star})^{\prime}(x)\leq 1/\sqrt{\ell_{V}}$. Hence, the final bound scales as
@@
$\|\bar{T}-\hat{T}\|_{L^{2}(\rho)}^{2}\leq{\varepsilon^{2}/\ell_{V}}\qquad\text{and}\qquad\|D(\bar{T}-\hat{T})\|_{L^{2}(\rho)}^{2}\leq{\varepsilon_{1}^{2}/\ell_{V}}\,.$ (23)
@@
We first make a series of reductions. By assumption, $D\bar{T}\succeq\alpha I$, and by definition, $\hat{T}$ is of the form $\alpha\,{\operatorname{id}}+\sum_{T\in\mathcal{M}}\lambda_{T}T+v$. By replacing $\bar{T}$ with $\bar{T}-\alpha\operatorname{id}$, it suffices to prove the following statement: assuming that $0\preceq D\bar{T}\preceq\ell_{V}^{-1/2}\,I$ together with the second derivative bound on $\bar{T}$, there exists $\hat{T}$ of the form $\sum_{T\in\mathcal{M}}\lambda_{T}T+v$ such that (23) holds. However, from the structure of $\mathcal{M}$, the problem now separates across the coordinates and it suffices to prove this statement with $d=1$ and $\varepsilon$ replaced with $\varepsilon/\sqrt{d}$.
Truncation procedure. We will construct $\hat{T}$ so that $\bar{T}(-R)=\hat{T}(-R)$ and $\bar{T}(+R)=\hat{T}(+R)$. Assuming that this holds, the bound on $\bar{T}^{\prime}$ and the fact that $\hat{T}$ is constant on $(-\infty,-R]$ and on $[+R,+\infty)$ readily imply
$|\bar{T}(x)-\hat{T}(x)|\leq\frac{1}{\sqrt{\ell_{V}}}\,(|x|-R)\,,\qquad\text{for}\leavevmode\nobreak\ |x|\geq R\,.$
@@
$\int_{\mathbb{R}\setminus(-R,+R)}|\bar{T}-\hat{T}|^{2}\,\mathrm{d}\rho\leq\frac{1}{\sqrt{2\uppi}\,\ell_{V}}\int_{\mathbb{R}\setminus(-R,+R)}{(|x|-R)}^{2}\exp(-x^{2}/2)\,\mathrm{d}x\,.$
@@
Standard Gaussian tail bounds and the Cauchy–Schwarz inequality imply that with the choice $R\asymp\sqrt{\log(1/(\ell_{V}\varepsilon^{2}))}$, we obtain $\|\bar{T}-\hat{T}\|_{L^{2}(\tilde{\rho})}^{2}\vee\|\bar{T}^{\prime}-\hat{T}^{\prime}\|_{L^{2}(\tilde{\rho})}^{2}\lesssim\varepsilon^{2}$, where $\tilde{\rho}$ is the Gaussian measure restricted to the set $\mathbb{R}\setminus[-R,R]$.
@@
With this choice of $\delta$, we then obtain $\varepsilon_{1}=\widetilde{O}(\sqrt{d}\,\kappa\delta)=\widetilde{O}(\kappa^{1/2}d^{1/4}\varepsilon^{1/2})$.
Size of the generating family. Finally, the size of $\mathcal{M}$ is $O({d}R/\delta)=\widetilde{O}(\kappa^{1/2}d^{{5}/4}/\varepsilon^{1/2})$, which completes the proof. ∎
@@
The key property of the piecewise cubics is that $(\psi^{\rm cub,\pm})^{\prime}(0)=(\psi^{\rm cub,\pm})^{\prime}(1)=0$. This means that even after adding the piecewise cubics, $\bar{T}^{\prime}$ and $\hat{T}^{\prime}$ agree at all of the endpoints of the sub-intervals. However, we must check that adding these piecewise cubics does not destroy the approximation rates (25) and (26). Since $|(\psi^{\rm cub,\pm}_{I})^{\prime}|\lesssim 1/\delta$, the bound on the coefficients for the piecewise cubics shows that the derivative of the piecewise cubic part of $\hat{T}$ is bounded in magnitude by $O(\kappa^{2}R^{2}\delta^{2}/\sqrt{\ell_{V}})$, so that (25) is intact. Similarly, (26) is also intact, either by integrating (25) or by using the bound on the coefficients of the piecewise cubics. Thus, $|\bar{T}(x)-\hat{T}(x)|\lesssim\kappa^{2}R^{2}\delta^{3}/\sqrt{\ell_{V}}$, and setting this to be at most $\varepsilon/\sqrt{d\ell_{V}}$ yields the choice $\delta=\widetilde{\Theta}(\varepsilon^{1/3}/(\kappa^{2/3}d^{1/6}))$.
Size of the generating family. The size of the generating family is then $O(dR/\delta)=\widetilde{O}(\kappa^{2/3}d^{7/6}/\varepsilon^{1/3})$, which completes the proof. ∎
@@
Proof of Lemma 5.13.
We restrict our attention to the piecewise linear family denoted $\mathcal{M}$ in dimension one with $|\mathcal{M}|=J$. This suffices due to the tensorization property of $\Upsilon$, see the remark after the definition of $\Upsilon$. It suffices to prove, for all $\lambda\in\mathbb{R}^{J}$,
$\bigl{\|}\sum_{T\in\mathcal{M}}\lambda_{T}T^{\prime}\bigr{\|}_{L^{2}(\rho)}^{2}\leq\Upsilon\,\bigl{\|}\sum_{T\in\mathcal{M}}\lambda_{T}T\bigr{\|}_{L^{2}(\rho)}^{2}\,,$
where $\rho=\mathcal{N}(0,1)$. We truncate the domain of $\rho$ to $[-R,R]$, where $R\asymp\sqrt{\log(1/(\ell_{V}\varepsilon^{2}))}$. On some interval $[a,a+\delta]$, note that
$T^{\lambda}(x)=T^{\lambda}(a)+\lambda_{T}\,((x-a)/\delta)_{+}\,,\quad DT^{\lambda}(x)=\lambda_{T}/\delta\,.$
It thus suffices to prove the statement on such an interval. This is equivalent to proving that
$\int_{a}^{a+\delta}\Bigl{(}\frac{\lambda_{T}}{\delta}\Bigr{)}^{2}\,\rho({\rm d}x)\leq\Upsilon\int_{a}^{a+\delta}\Bigl{(}T^{\lambda}(a)+\lambda_{T}\,\frac{x-a}{\delta}\Bigr{)}^{2}\,\rho({\rm d}x)\,.$
Rearranging, it suffices to show that
$\delta^{-2}\Upsilon^{-1}\leq\frac{\inf_{m\in\mathbb{R}}\int_{a}^{a+\delta}\bigl{(}\frac{x-a}{\delta}-m\bigr{)}^{2}\,\rho({\rm d}x)}{\int_{a}^{a+\delta}\rho({\rm d}x)}=\frac{{\rm{var}}X}{\delta^{2}}\,,$
or $\Upsilon^{-1}\leq{\rm{var}}X$, with $X\sim\rho|_{[a,a+\delta]}$.
Letting $m_{a}\coloneqq\mathbb{E}X$, suppose WLOG $m_{a}\leq a+\delta/2$. We compute
[…]
provided $\delta\lesssim 1/R$; indeed, for this choice of $\delta$, $|{\log\rho(x)-\log\rho(y)}|\lesssim 1$ for all $x,y\in[a,a+\delta]$. Stringing together the inequalities, we obtain the desired claim. ∎
@@
Recall our definition of $\pi^{\star}$ with components $\pi_{i}\propto\exp(-V_{i})$ with
$V_{i}(x_{i})\int_{\mathbb{R}^{d-1}}V(x_{1},\ldots,x_{d})\,\bigotimes_{j\neq i}\pi^{\star}_{j}(\!\,\mathrm{d}x_{j})$
Assuming the first claim, we can prove the second by applying the Brascamp–Lieb inequality (Brascamp and Lieb,, 1976):
@@
Finally, we also prove the bound on $\Xi$ for the piecewise linear dictionary.
Proof of Lemma 5.15.
If we can show that $Q\succeq\gamma I$ for some $\gamma>0$, then
$\langle Q^{-1},\bar{Q}(x)\rangle$ $\leq\gamma^{-1}\operatorname{tr}\bar{Q}(x)=\gamma^{-1}\sum_{T\in\mathcal{M}}{T(x)}^{2}\leq\gamma^{-1}J\,,$
where we use the fact that the elements of the piecewise linear dictionary are uniformly bounded by $1$.
To prove the lower bound on $Q$, we note that for any $\lambda\in\mathbb{R}^{J}$,
$\langle\lambda,Q\,\lambda\rangle$ $=\Bigl{\lVert}\sum_{T\in\mathcal{M}}\lambda_{T}T\Bigr{\rVert}_{L^{2}(\rho)}^{2}\,.$
On an interval $[a,a+\delta]$, since $T^{\lambda}(x)=T^{\lambda}(a)+\lambda_{T}\,((x-a)/\delta)_{+}$,
$\int_{a}^{a+\delta}T^{\lambda}(x)^{2}\,\rho(\mathrm{d}x)$ $=\int_{a}^{a+\delta}\Bigl{(}T^{\lambda}(a)+\lambda_{T}\,\frac{x-a}{\delta}\Bigr{)}^{2}\,\rho(\mathrm{d}x)\geq\lambda_{T}^{2}\inf_{m\in\mathbb{R}}\int_{a}^{b}\bigl{(}\frac{x-a}{\delta}-m\bigr{)}^{2}\,\rho(\mathrm{d}x)$
$\gtrsim\lambda_{T}^{2}\delta^{2}\int_{a}^{a+\delta}\rho(\mathrm{d}x)\,,$
where we used the variance bound from the proof of Lemma 5.13. Summing across the intervals, we find that $\langle\lambda,Q\,\lambda\rangle\gtrsim\delta^{2}\sum_{T\in\mathcal{M}}\lambda_{T}^{2}$, so we can take $\gamma\asymp\delta^{2}$. This leads to an upper bound on $\Xi$ of order $\delta^{-2}\asymp J^{2}$. ∎
Appendix D Remaining implementation details
D.1 Product Gaussian mixture
Let $V_{1}$ (resp. $V_{2}$) be the potential for a univariate Gaussian mixture with weights $w_{1,1}$ and $w_{1,2}$ (resp. $w_{2,1}$ and $w_{2,2}$) that sum to unity, and centers $m_{1,1}$ and $m_{1,2}$ (resp. $m_{2,1}$ and $m_{2,2}$), where all the mixture components have unit variance. Then, $V:\mathbb{R}^{2}\to\mathbb{R}$ defined by $V(x,y)=V_{1}(x)+V_{2}(y)$ is the potential for the Gaussian mixture with mean-weight pairs given by
[…]
$([m_{1,2},m_{2,1}],w_{1,2}w_{2,1}),([m_{1,2},m_{2,2}],w_{1,2}w_{2,2})\}\,.$
We take $m_{1,1}=m_{2,1}=2$, $m_{1,2}=m_{2,2}=-2$, with $w_{1,1}=w_{2,2}=0.25$ and $w_{1,2}=w_{2,1}=0.75$. As for the hyperparameters of our model, we chose $J=28$, $\alpha=0.1$, a step-size $h=10^{-3}$ (for both $\lambda$ and $v$), ran for 3000 iterations, and initialized at $\lambda^{(0)}=0_{2\times J}\in\mathbb{R}^{2\times J}$, and $v^{(0)}=0_{2}\in\mathbb{R}^{2}$. The KDE plots were generated via sklearn, after we generated 50,000 samples from the ground truth density and from our algorithm.
D.2 Non-isotropic Gaussian
We generated $A\in\mathbb{R}^{d\times d}$ with entries $A_{i,j}\sim\mathcal{N}(0,1)$, and defined $\Sigma=AA^{\top}$ for $d=5$, which is fixed once and for all. We computed the optimal $\alpha^{*}=1/\sqrt{L_{V}}$, since the potential is a Gaussian. For the remaining hyper-parameters of our model, we chose $J=28$, a step-size $h=10^{-4}$ (for both $\lambda$ and $v$), ran for $2000$ iterations, and initialized at $\lambda^{(0)}=\bm{1}_{d\times J}\in\mathbb{R}^{d\times J}$, the all-ones matrix, and $v^{(0)}=0_{d}\in\mathbb{R}^{d}$. At each step, we computed $\hat{\Sigma}_{\text{MF}}$ by pushing forward 10,000 samples, computing the empirical covariance, and computing the Bures–Wasserstein distance to $\Sigma_{\text{MF}}$.
We now compute the fact that $\Sigma_{\text{MF}}$ is diagonal with components $1/(\Sigma^{-1})_{i,i}$ for $i\in[d]$. Recall the KL divergence between two Gaussians with mean zero is given by
$\operatorname*{KL}(\mathcal{N}(0,A)\!\;\|\;\!\mathcal{N}(0,\Sigma))=\frac{1}{2}\bigl{[}\operatorname{tr}(\Sigma^{-1}A)-d+\log\det(\Sigma^{-1})-\log\det(A)\bigr{]}\,.$
Now, we impose that $A$ is a diagonal matrix with entries $A_{i,i}=a_{i}$ for some $a_{i}\geq 0$. In this case, up to constants denoted by $C$, the above reads
$\operatorname*{KL}(\mathcal{N}(0,A)\!\;\|\;\!\mathcal{N}(0,\Sigma))=\frac{1}{2}\sum_{i=1}^{d}\bigl{[}(\Sigma^{-1})_{i,i}a_{i}-\log(a_{i})\bigr{]}+C\,.$
Taking the derivative in $a_{i}$, we see that the optimality conditions yield
$1/(\Sigma^{-1})_{i,i}=a_{i}^{\star}$
for every $i\in[d]$, which completes the calculation.
D.3 Bayesian logistic regression
We first randomly drew $\theta^{\star}\sim\mathcal{N}(0,I_{d})$ in $d=20$ as the ground truth parameter. Further, we let $n=100$ and randomly generated $X\in\mathbb{R}^{n\times d}$ as in the non-isotropic Gaussian experiment (here, $X$ takes the role of $A$), but we divided the matrix by $\lambda_{\max}(X^{\top}X)$ for normalization purposes. Subsequently, $Y_{i}$ was generated for each $i$ independently according to
$Y_{i}\mid X_{i}\sim\text{Bern}(\exp(\theta^{\top}X_{i}))\,,$
where $X_{i}$ is a row of $X$. Using this data, and assuming an improper (Lebesgue) prior on $\theta$, the potential of the posterior is given by
$V(\theta)=\sum_{i=1}^{n}\bigl{[}\log(1+\exp(\theta^{\top}X_{i}))-Y_{i}\,\theta^{\top}X_{i}\bigr{]}\,.$
With access to $V$ and $\nabla V$, we ran standard Langevin Monte Carlo (LMC) for 5000 iterations with a step size of $h=10^{-2}$, where we generated 2000 samples.
For the hyperparameters of our model, we chose $J=28$, $\alpha=0.1$, a step size $h=10^{-2}$ for the $\lambda$ iterates, and $h_{v}=10^{-1}$ for updating $v$, and ran for $2000$ iterations. We initialized at $\lambda^{(0)}=\bm{1}_{d\times J}/(Jd)\in\mathbb{R}^{d\times J}$, and $v^{(0)}=0_{d}\in\mathbb{R}^{d}$. The final histograms were generated using 2000 samples from both the mean-field VI algorithm and LMC.
Appendix E Proofs for Section 7
In this section, we derive the gradient flows in Section 7.
Proof of Theorem 7.1.
We refer to Lambert et al., (2022, Appendix F) for the relevant background. The first variation of the functional $\mathcal{F}(P)\coloneqq\operatorname*{KL}(\mu_{P}\!\;\|\;\!\pi)$ is given by
$\delta\mathcal{F}(P):(\lambda,v)\mapsto\int(V+\log\mu_{P}+1)\,\mathrm{d}\mu_{\lambda,v}=\int\log\frac{\mu_{P}}{\pi}\,\mathrm{d}\mu_{\lambda,v}+1\,.$ (29)
Therefore, the Wasserstein gradient is given by
$\nabla_{\mathbb{W}}\mathcal{F}(P)(\lambda,v)$ $=\Bigl{(}Q^{-1}\,\nabla_{\lambda}\int\log\frac{\mu_{P}}{\pi}\,\mathrm{d}\mu_{\lambda,v},\,\nabla_{v}\int\log\frac{\mu_{P}}{\pi}\,\mathrm{d}\mu_{\lambda,v}\Bigr{)}\,.$ (30)
These terms are further computed as follows. First,
$\partial_{\lambda_{T}}\int\log\frac{\mu_{P}}{\pi}\,\mathrm{d}\mu_{\lambda,v}$ $=\partial_{\lambda_{T}}\int\log\frac{\mu_{P}}{\pi}\circ T^{\lambda,v}\,\mathrm{d}\rho=\int\bigl{\langle}\nabla\log\frac{\mu_{P}}{\pi}\circ T^{\lambda,v},T\bigr{\rangle}\,\mathrm{d}\rho\,.$
Similarly, we have
$\nabla_{v}\int\log\frac{\mu_{P}}{\pi}\,\mathrm{d}\mu_{\lambda,v}$ $=\int\nabla\log\frac{\mu_{P}}{\pi}\circ T^{\lambda,v}\,\mathrm{d}\rho\,.$
This concludes the proof. ∎
Proof of Theorem 7.2.
This theorem follows from the expression of the first variation computed in (29), see Lambert et al., (2022, Appendix H). ∎